Over the last three evenings, Jane received a total of 68 phone calls at the call center. The third evening, she received 3 times as many calls as the first evening. The second evening, she received 7 fewer calls than the first evening. How many phone calls did she receive each evening?
Answer by ikleyn(52848)

.

    x + (x-7) + 3x = 68.

    5x - 7 = 68

    5x     = 68 + 7 = 75

     x              = 75/5 = 15 phone calls.


ANSWER.  15 phone calls first evening;  15-7 = 8 phone calls seconf evening and  3*15 = 45 phone calls third evening.

Solved.

You may check the answer by adding all 3 numbers in it.